Find the solutions to the system of equations by the addition method. x-5y=-47 -4x-5y=-37 is the answer (-1,9)
Answer by drk(1908)

Here are the two equations:
(i) 1x - 5y = -47
(ii) -4x - 5y = -37
multiply (i) by -1 to get
(iii) -1x + 5y = +47
(ii) -4x - 5y = -37
add down and find x.
(iv) -5x = 10
x = -2
solve for y.
y = 9.
(-2,9)